OpenStudy (anonymous):

The integral of 3sin^4(3x)

OpenStudy (anonymous):

\[\int\limits 3\sin ^{4} (3x)\]

OpenStudy (anonymous):

dont worry babes muku is here ;)

OpenStudy (anonymous):

\(\sin^4(3x)=\sin^2(3x)\sin^2(3x)=\sin^2(3x) (1-\cos^2(3x))=\sin^2(3x)-\sin^2(3x)\cos^2(3x)\\=\sin^2(3x)-\frac{1}{4} \sin^2(6x) \)

OpenStudy (experimentx):